OpenStudy (anonymous):

Use L'Hoptals rule: lim x->0 (cos(x))^(1/x^2) I have no idea where to start since this isn't a fraction

OpenStudy (zarkon):

make it into a fraction by using logs

OpenStudy (anonymous):

As @Zarkon said \[ \ln \left(\sqrt[x^2]{\cos (x)}\right)=\frac{\ln (\cos (x))}{x^2} \]

OpenStudy (anonymous):

if x->0 we have 0/0 so you can apply L'Hopital's rule

OpenStudy (anonymous):

Take the derivative of the numerator over the derivative of the denominator, you get after simplification \[ -\frac{\sin (x)}{(2 x) \cos (x)}\to -\frac 1 2 \] When x goes to zero. It is easy now to find the original limit

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

YW. Do you know how to finish it?

OpenStudy (anonymous):

Yeah I at first got e^0.5 but then your work showed me that Dx(cosx) was -sinx/cosx not sinx/cosx. So then I got e^-0.5

OpenStudy (anonymous):

Yes, the answer is \[ \frac{1}{\sqrt{e}} \]

OpenStudy (anonymous):

Thank you I knew how to start it after Zarkon's reply but I give eliassaab the best response because of my error taking the derivative that he corrected